# Exploring the Physics of the Plasma Liner Experiment: A Multi-dimensional Study with FLASH, OSIRIS, and HELIOS

### v1

@@ Abstract
The Plasma Liner Experiment (PLX) at Los Alamos National Laboratory (LANL) is a platform that seeks to achieve fusion via a concept known as Plasma-Jet-Driven Magneto-Inertial Fusion (PJMIF). The experiment consists of three main phases: (1) target formation in which up to four plasma guns shoot magnetized hydrogen or deuterium-tritium jets to form a quasi-spherical target, (2) liner formation in which a constellation of 36 guns fire high-atomic-number (e.g., xenon) jets to form a liner shell, and (3) target compression in which the formed liner implodes the pre-formed target. Each phase of the PLX probes different plasma regimes with different physics at play, thus we simulated each phase separately and with multiple codes. Here we highlight some of the 1D, 2D, and 3D simulation results of all three phases from the FLASH, OSIRIS, and HELIOS codes. Some of the key physical processes involved include shock dynamics, kinetic effects, anisotropic thermal conduction, resistive magnetic diffusion, radiation transport, and magnetized jet dynamics. Our simulations show that the PLX can form a preheated (40 eV), magnetized (electron Hall parameter 1) target plasma, and a quasi-collisional liner shell that can subsequently compress the target to fusion-relevant conditions (e.g., temperatures 1 keV).
@@ I Introduction
The pursuit of practical fusion energy remains a grand scientific and engineering challenge, with several promising approaches under active investigation Ikeda (2007); Buttery et al. (2019); Zylstra et al. (2022); Abu-Shawareb et al. (2024). Among them, Plasma-Jet-Driven Magneto-Inertial Fusion (PJMIF) represents a hybrid method that seeks to combine the advantages of both magnetic and inertial confinement fusion Hsu et al. (2012); Thio et al. (2019). This novel concept employs high-velocity plasma jets to form a spherically symmetric liner, which compresses a magnetized fuel target to fusion conditions. The PJMIF approach has the potential to achieve high energy gain at relatively low driver energies, making it an attractive alternative to conventional fusion concepts.
@@
The Plasma Liner Experiment (PLX) at Los Alamos National Laboratory (LANL) serves as a critical testbed for investigating the feasibility of PJMIF Merritt et al. (2014); Langendorf et al. (2018); LaJoie et al. (2024). The heavy, fast-moving plasma liner used to drive the implosion of the plasma target is formed from the merging and spherical convergence of an array of 36 supersonic plasma jets traveling at km/s Yates et al. (2020). Recent advancements have improved our understanding of PLX processes such as liner uniformity and jet merging dynamics and have guided the development of optimized jet configurations and injection parameters to enhance liner performance LaJoie et al. (2023, 2024). A compatible magnetized plasma target that takes advantage of the high implosion speed of a spherically imploding plasma liner is injected into the vessel by merging 46 magnetized plasma jets with a magnetic field strength of 1000 G at the gun nozzles. The collision of these plasma jets is expected to induce turbulence in the plasma target, resulting in tangled fields, which can provide very long connection lengths between the core and the liner surface thereby reducing thermal conduction losses from the fuel plasma to the colder liner Ryutov (2009); Hsu and Langendorf (2019a); Chu et al. (2023a). A comprehensive investigation into plasma target formation on PLX is scheduled to commence in late 2025.
@@
The conditions in PJMIF implosions span a wide range of plasma parameter space and physical regimes, ranging from kinetic to fluid collisionality, optically thin to thick radiation transport, and weak to moderate magnetization. Some of the important physical processes include cold, high velocity jets merging at oblique angles during liner formation, target formation of a pre-heated magnetized plasma, anisotropic magnetized thermal conduction, and ultimately fusion thermonuclear burn and charged particle transport Chu et al. (2023b). Due to the wide range of relevant physics, previous studies in PJMIF have often employed analytical or semi-analytical approaches to predict performance. Langendorf and Hsu (2017); Hsu and Langendorf (2019b) Integrated simulation efforts have been attempted and have addressed many first-order questions Awe et al. (2011); Davis et al. (2012); Thoma et al. (2013); Schillo and Cassibry (2020); Thompson and Cassibry (2020), but have often been limited in the breadth of the included physical effects. In this work, an attempt is made to rigorously simulate as many aspects of the concept as feasible with leading modern computational capabilities, and give improved insight into the detailed performance of PJMIF conceptual designs.
@@
The structure of the paper is as follows: in Section II we describe the three codes used in this work: FLASH, OSIRIS and HELIOS. Then the paper is divided by the three phases of PLX: target formation (Section III), liner formation (Section IV), and target compression (Section V). Lastly, we summarize our conclusions in Section VI.
@@ II Numerical Methods
For the radiation-MHD modeling of the various stages of the PLX platform we employ the FLASH code Fryxell et al. (2000). FLASH is a publicly-available, parallel, multi-physics, adaptive mesh refinement (AMR), finite-volume Eulerian hydrodynamics and MHD code, developed at the University of Rochester by the Flash Center for Computational Science (for more information on the FLASH code, visit: https://flash.rochester.edu). FLASH scales well to over 100,000 processors in modern high-performance computing architectures, and uses a variety of parallelization techniques like domain decomposition, mesh replication, and threading, to optimally utilize hardware resources. The FLASH code has a world-wide user base of more than 4,600 scientists, and more than 1,300 papers have been published using the code to model problems in plasma astrophysics, combustion, computational fluid dynamics, high energy density physics (HEDP), and fusion energy research.
@@
Over the past decade and under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration (NNSA), the Flash Center has added in FLASH extensive HEDP and extended-MHD capabilities Tzeferacos et al. (2015) that make it an ideal tool for the multi-physics modeling of the PLX platform. These include state-of-the art hydrodynamic and MHD shock-capturing solvers, Lee (2013) extended to a three-temperature formalism, Tzeferacos et al. (2015) anisotropic thermal conduction that utilizes high-fidelity magnetized heat transport coefficients, Ji and Held (2013) electron-ion heat exchange, multi-group radiation diffusion4, tabulated multi-material equations of state (EOS) and opacities, and numerous synthetic diagnostics Tzeferacos et al. (2017). FLASH’s newest algorithmic developments also include a complete generalized Ohm’s law that incorporates all extended MHD terms of the Braginskii formulation. Braginskii (1965) The new extended MHD capabilities are integrated with state-of-the-art transport coefficients, Davies et al. (2021) developed with support from the U.S. DOE Advanced Research Projects Agency–Energy (ARPA-E) BETHE program. The FLASH code and its capabilities for magnetized HEDP have been verified through several benchmarks and code-to-code comparisons, Fatenejad et al. (2013); Orban et al. (2013, 2022); Sauppe et al. (2023) as well as through direct application to numerous plasma physics experiments, Meinecke et al. (2014, 2015); Li et al. (2016); Tzeferacos et al. (2018); Rigby et al. (2018); White et al. (2019); Chen et al. (2020); Bott et al. (2021); Meinecke et al. (2022) leading to innovative science and publications in high-impact journals.
@@
For the kinetic modeling of target and liner plasma jet merging processes we employ the OSIRIS code.Fonseca et al. (2002); Hemker et al. (1999) OSIRIS is a state-of-the-art, fully explicit, multi-dimensional (1D, 2D slab and -, and 3D), fully parallelized, fully relativistic, particle-in-cell (PIC) code. OSIRIS is jointly developed by Instituto Superior Técnico Lisboa and University of California, Los Angeles (for more information on the OSIRIS code, visit: https://osiris-code.github.io). OSIRIS is highly optimized on a single core, and it scales well to over 1.5 million cores on the Sequoia supercomputer at Lawrence Livermore National Laboratory, maintaining 97% and 78% efficiency on weak and strong scaling, respectively. It has been run on many of the DOE and NSF leadership class computing facilities for over 15 years. On Blue Waters it has achieved 2.2 Pflops on a full-machine benchmark using more than 10 trillion particles. The OSIRIS code is used world-wide by more than 25 user groups to model problems in laser-plasma interaction, wake field acceleration, ion acceleration, plasma astrophysics such as shock formation, and fusion energy research.
@@
OSIRIS includes higher order particle shapes, which limits self-heating effects related to aliasing, and includes energy and momentum conserving algorithms. The parallelization is done either using domain decomposition with MPI only or by using a hybrid approach where MPI is used across nodes and OpenMP parallelization is used within a node; both approaches support dynamic load balancing at the MPI level.Miller et al. (2021) The dynamic load balancing feature can greatly speed up the plasma jet collision simulations because the computation time for each time step is roughly proportional to the number of particles inside the MPI nodes. The spatial distribution of the number of particles changes drastically as the simulations progress. There is a relativistic two-body Coulomb collision model Nanbu (1997); Takizuka and Abe (1977); Peano et al. (2009) as well as a hybrid algorithm Fiuza et al. (2011) which pushes particles throughout all regions, but in a high-density region (collisional), MHD field equations are used to solve for the fields. This is useful in simulations of plasmas with a large density range where some parts of the system are collisionless while others are collisional.
@@
Results are also compared with HELIOS, a lightweight 1D Lagrangian fluid code that includes physics packages useful for HEDP such as tabulated EOS, flux-limited radiation diffusion, and thermal conduction. MacFarlane et al. (2006) Of particular interest in this study, the 1D Lagrangian formulation will exactly preserve a material interface (between liner and target) with no numerical diffusion or mixing between the two species. In a realistic case, mixing may occur due to various physical mechanisms including Rayleigh-Taylor instabilities, but the comparison of the 1D Lagrangian-to-Eulerian codes allows a convenient way to gauge the pure numerical diffusion that may be occurring in the Eulerian codes as a function of grid resolution, separate from physical mechanisms.
@@ III Target Formation
We ran a 2D axisymmetric FLASH simulation of two counter-propagating jets to demonstrate how a target plasma is formed from the jet collision. The code solves the three-temperature resistive MHD equations with anisotropic thermal conduction, electron-ion heat exchange, and multi-group radiation (six groups). For the jets, we used newer thermal conductivity Ji and Held (2013) and magnetic resistivity Davies et al. (2021) coefficients, and PROPACEOS tables for multi-group opacities. The domain extends axially from 137 cm to -137 cm (the diameter of the PLX chamber), and spans 68.5 cm radially. We used adaptive mesh refinement to achieve a maximum effective resolution of nearly 32 cells across the initial outer jet radius.
@@
The jets are modeled as fully-ionized hydrogen with an ideal EOS (). They are initialized with an inner radius of 1.5 cm (the plasma gun wire radius) and an outer radius of 4.25 cm. Between these radii, the jets are given a mass density of g/cm3, a velocity of 80 km/s, electron and ion temperatures of 1 eV, and a radiation temperature of 0.1 eV. These parameters are enforced as inflow conditions at the lower and upper domain boundaries, which are also treated as solid reflecting walls from to 1.5 cm and beyond 4.25 cm. The solid wall regions of the domain boundaries do not allow any hydrodynamic motion, but they will allow thermal, radiation, and magnetic diffusion to pass through them. A radially-varying azimuthal magnetic field is imposed for the entire radial extent at these boundaries: it has a peak value of 2500 G at 1.5 cm with a linear profile from cm to 1.5 cm, a profile from 1.5 cm to 4.25 cm, and a value of zero beyond 4.25 cm. This magnetic field profile is also imposed every time step in the first row of cells inside the domain; this helps reduce the imbalance of hydromagnetic pressure with the inflow conditions for better control of the jet velocity.
@@
Since FLASH cannot simulate a true vacuum, the PLX chamber or ambient region is initialized with a low density of g/cm3, and with thermal conduction and radiation absorption and emission turned off. The electron and ion temperatures are initialized to 1 eV and radiation temperature to 0.1 eV everywhere inside the domain. The difference in density between the vacuum and jet material creates a pressure imbalance which increases the forward jet velocity and also causes the jets to expand radially into the vacuum. There are also forces from the azimuthal magnetic field that pinch the jet towards during its propagation. Once the pinched jet material reaches the symmetry axis, it has no choice but to move axially, which leads to a streamlining of the jet head, observed as an overall acceleration of the jet; this is a well-documented phenomenon sometimes referred to as the nose-cone effect (see Hansen et al. (2015) and references therein). The combined effects of the enforced jet inflow velocity, pressure imbalance, and pinch forces result in an average jet velocity (from s to collision) of about 125 km/s, which is higher than the reported experimental values Hsu et al. (2015) but not necessarily inconsistent. Previous experiments measured the jet velocities at positions near the gun and slightly downstream and reported a possible range of 30-100 km/s. Given the aforementioned acceleration mechanisms, an average value of 125 km/s over the entire axial extent is plausible.
@@
Fig. 1 shows a density and electron temperature map of the jets after they have collided and formed a target plasma. Note that some amount of jet material is left behind or expands outwards away from the target-forming region. This effect can be mitigated by adding more jets; the experimental platform currently employs four target-forming jets, but may use more in the future. Also, some of the jet material will be swept up by the liner-forming jets during the subsequent liner-forming process. As the jets collide, material is forced outwards radially created an elongated pancake-like target. This effect can also be mitigated by adding more jets to the process, which would result in a more spherical target. In the end, the target reaches a volume-averaged density of approximately g cm-3.
@@
The temperature map of Fig. 1 shows that the hottest plasma forms a shell around the newly formed target. These outermost regions of the target plasma are similar to accretion shocks where jet material is piling up and ramming into the expanding target. Since the magnetic field is azimuthal and the hot shell has a relatively lower magnetic resistivity, the magnetic field has increased difficulty leaving the shell region. This in turn can assist in thermally insulating the target via reduced thermal conductivity, which is included in the anisotropic thermal conduction implementation in FLASH. Some amount of preheat before the incoming liner compresses the target is crucial to the overall process, and we observe peak temperatures of around 40 eV in the jet collision simulation.
@@
Our analysis included volume-averaging several quantities over a region about the center of the domain from to cm and to cm. In this region, the target attains a volume-averaged number density of approximately cm-3, which would increase experimentally with more jets. We also expect the final volume-averaged magnetic field strength, 24.4 G, to increase with the number of jets. Fig. 2 shows the volume-averaged Hall parameters for electrons () and for ions () as a function of time. The target plasma reaches , which is consistent with previous FLASH and OSIRIS simulations of target formation Wen et al. (2022) that reported electron Hall parameters greater than unity. Also, the attained average plasma (ratio of thermal pressure to magnetic pressure) was slightly greater than 100. This indicates that the magnetic field is strong enough to assist in thermally insulating the target plasma while not being dynamically significant, which is one of the primary goals of the PLX target formation process.
@@
The results from this jet collision simulation are used to inform the initial conditions for 3D experimental-scale target compression simulations presented later in this work (Section V.2).

### v2

@@ Abstract
The Plasma Liner Experiment (PLX) at Los Alamos National Laboratory (LANL) is a platform that seeks to achieve fusion via the Plasma-Jet-Driven Magneto-Inertial Fusion (PJMIF) concept. The experiment utilizes a constellation of plasma guns to generate fusion-relevant conditions and consists of three main phases: (1) target formation, in which up to four plasma guns shoot magnetized hydrogen or deuterium-tritium jets to form a quasi-spherical target, (2) liner formation, in which a 36 guns fire high-atomic-number (e.g., xenon) jets to form a liner shell, and (3) target compression, in which the formed liner implodes the pre-formed target. Each phase of the PLX operates in different plasma regimes, with different physics at play, thus each phase must be simulated separately with appropriate codes. In this study we highlight 1-, 2-, and 3-D simulation results of all three phases using the FLASH, OSIRIS, and HELIOS codes. Some of the key physical processes involved include shock dynamics, kinetic effects, anisotropic thermal conduction, resistive magnetic diffusion, radiation transport, and magnetized jet dynamics. The simulations show that the PLX can form a preheated (40 eV), magnetized (electron Hall parameter 1) target plasma, and a quasi-collisional liner shell, which can subsequently compress the target to fusion-relevant conditions, reaching temperatures in excess of 1 keV.
@@ I Introduction
The pursuit of practical fusion energy remains a grand scientific and engineering challenge, with several promising approaches under active investigation ikeda2007progress; buttery2019diii; zylstra2022burning; abu2024achievement. Among them, Plasma-Jet-Driven Magneto-Inertial Fusion (PJMIF) represents a hybrid method that seeks to combine the advantages of both magnetic and inertial confinement fusion Hsu2012; Thio2019. This novel concept employs high-velocity plasma jets to form a spherically symmetric liner, which compresses a magnetized fuel target to fusion conditions. The PJMIF approach has the potential to achieve high energy gain at relatively low driver energies, making it an attractive alternative to conventional fusion concepts.
@@
The Plasma Liner Experiment (PLX) at Los Alamos National Laboratory (LANL) serves as a critical testbed for investigating the feasibility of PJMIF merritt2014experimental; langendorf2018experimental; lajoie2024formation. The heavy, fast-moving plasma liner used to drive the implosion of the plasma target is formed from the merging and spherical convergence of an array of 36 supersonic plasma jets traveling at km/s yates2020experimental. Recent advancements have improved our understanding of the PLX properties, such as liner uniformity and jet merging dynamics, and have guided the development of novel, optimized jet configurations and injection parameters to enhance liner performance lajoie2023multi; lajoie2024formation. A compatible magnetized plasma target that takes advantage of the high implosion speed of a spherically imploding plasma liner is injected into the vessel by merging magnetized plasma jets, whose magnetic field strength is 1 kG at the gun nozzles. The collision of these plasma jets is expected to induce turbulence in the plasma target, resulting in tangled fields, which can in turn provide very long conduction lengths between the core and the liner surface thereby reducing thermal conduction losses from the fuel plasma to the colder liner ryutov2009adiabatic; hsu2019magnetized; chu2023characterization.
@@
The conditions in PJMIF implosions span a wide range of plasma parameter space and physical regimes, ranging from kinetic to fluid collisionality, optically thin to thick radiation transport, and weak to moderate magnetization. Some of the pertinent physical processes include cold, high-velocity jets merging at oblique angles during liner formation, target formation of a preheated magnetized plasma, anisotropic magnetized thermal conduction, and ultimately fusion thermonuclear burn and charged particle transport chu2023experimental. Due to the wide range of relevant physics, previous studies in PJMIF have often employed analytical or semi-analytical approaches to predict performance. Langendorf2017; Hsu2019 Integrated simulation efforts have been attempted and have addressed many first-order questions Awe2011; Davis2012; Thoma2013; Schillo2020; Thompson2020, but have often been limited in the breadth of the physical effects they could include. In this work, an attempt is made to rigorously simulate as many aspects of the concept as feasible with leading modern computational capabilities, and give improved insight into the detailed performance of PJMIF conceptual designs.
@@
The structure of the paper is as follows: in Section II we describe the three codes used in this work: FLASH, OSIRIS and HELIOS. Then, the paper is divided by the three phases of the PLX: target formation (Section III), liner formation (Section IV), and target compression (Section V). Lastly, we summarize our conclusions in Section VI.
@@ II Numerical Methods
For the radiation magneto-hydrodynamic (MHD) modeling of the various stages of the PLX platform we employ the FLASH code Fryxell2000. FLASH is a publicly-available, parallel, multi-physics, adaptive mesh refinement (AMR), finite-volume Eulerian hydrodynamics and MHD code, developed at the University of Rochester by the Flash Center for Computational Science (for more information on the FLASH code, visit: https://flash.rochester.edu). FLASH scales well to over 100,000 processors in modern high-performance computing architectures, and uses a variety of parallelization techniques like domain decomposition, mesh replication, and threading, to optimally utilize hardware resources. The FLASH code has a world-wide user base of more than 5,000 scientists, and more than 1,300 papers have been published using the code to model problems in plasma astrophysics, combustion, computational fluid dynamics, high energy density physics (HEDP), and fusion energy research.
@@
Over the past decade and under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration (NNSA), the Flash Center has added in FLASH extensive HEDP and extended-MHD capabilities Tzeferacos2015 that make it an ideal tool for the multi-physics modeling of the PLX platform. These include state-of-the art hydrodynamic and MHD shock-capturing solvers, Lee2013 extended to a three-temperature formalism, Tzeferacos2015 anisotropic thermal conduction that utilizes high-fidelity magnetized heat transport coefficients, JiHeld2013 electron-ion heat exchange, multi-group radiation diffusion, tabulated multi-material equations of state (EOS) and opacities, and numerous synthetic diagnostics Tzeferacos2017. FLASH’s newest algorithmic developments also include a complete generalized Ohm’s law that incorporates all extended-MHD terms of the Braginskii formulation. Braginskii1965 The new extended MHD capabilities are integrated with state-of-the-art transport coefficients, Davies2021 developed with support from the U.S. DOE Advanced Research Projects Agency–Energy (ARPA-E) BETHE program. The FLASH code and its capabilities for magnetized HEDP have been verified through several benchmarks and code-to-code comparisons, Fatenejad2013; Orban2013; Orban2022; Sauppe2023 as well as through direct application to numerous plasma physics experiments, Meinecke2014; Meinecke2015; Li2016; Tzeferacos2018; Rigby2018; White2019; Chen2020; Bott2021; Meinecke2022 leading to innovative science and publications in high-impact journals.
@@
For the kinetic modeling of target and liner plasma-jet merging processes we employ the OSIRIS code.Fonseca02; hemker00 OSIRIS is a state-of-the-art, fully explicit, multi-dimensional (1D, 2D slab and -, and 3D), fully parallelized, fully relativistic, particle-in-cell (PIC) code. OSIRIS is jointly developed by Instituto Superior Técnico Lisboa and the University of California, Los Angeles (for more information on the OSIRIS code, visit: https://osiris-code.github.io). OSIRIS is highly optimized on a single core, and it scales well to over 1.5 million cores on the Sequoia supercomputer at Lawrence Livermore National Laboratory, maintaining 97% and 78% efficiency on weak and strong scaling, respectively. It has been run on several of the U.S. DOE and National Science Foundation (NSF) leadership-class computing facilities for over 15 years. On Blue Waters it has achieved 2.2 Pflops on a full-machine benchmark using more than 10 trillion particles. The OSIRIS code is used world-wide by more than 25 user groups to model problems in laser-plasma interaction, wake field acceleration, ion acceleration, plasma astrophysics such as shock formation, and fusion energy research.
@@
OSIRIS includes higher order particle shapes, which limits self-heating effects related to aliasing, and includes energy- and momentum-conserving algorithms. The parallelization is done either using domain decomposition with MPI only or by using a hybrid approach where MPI is used across nodes and OpenMP parallelization is used within a node; both approaches support dynamic load balancing at the MPI level.Miller2021 The dynamic load balancing feature can greatly speed up the plasma-jet collision simulations because the compute time for each time step is roughly proportional to the number of particles in the MPI nodes. The spatial distribution of the number of particles changes drastically as the simulations progress. There is a relativistic two-body Coulomb collision model nanbu1997; takizuka1977; peano2009 as well as a hybrid algorithm Fiuza2011 which pushes particles throughout all regions, but in a high-density region (i.e., collisional), MHD field equations are used to solve for the fields. This is useful in simulations of plasmas with a large density range, where some parts of the system are collisionless while others are collisional.
@@
Results are also compared with HELIOS, a lightweight 1D Lagrangian fluid code that includes physics packages useful for HEDP such as tabulated EOS, flux-limited radiation diffusion, and thermal conduction. MacFarlane2006 Of particular interest in this study, the 1D Lagrangian formulation exactly preserves material interfaces (e.g., between the liner and the target) with no numerical diffusion or mixing between species. In a realistic case, mixing may occur due to various physical mechanisms, such as the Rayleigh-Taylor instability, but the comparison of 1D Lagrangian-to-Eulerian codes allows for a convenient way to gauge the impact of numerical diffusion in the Eulerian solve, as a function of grid resolution, separate from physical processes.
@@ III Target Formation
To model the targer formation phase of the PLX, we ran a 2D axisymmetric FLASH simulation of two counter-propagating jets to demonstrate how the target plasma is formed from the jet collision. The code solves the three-temperature resistive MHD equations with anisotropic thermal conduction, electron-ion heat exchange, and multi-group radiation (with six energy groups). For the jets, the new implementations of thermal conductivityJiHeld2013 and magnetic resistivityDavies2021 coefficients are used, coupled with PROPACEOS tables for the frequency-dependent opacity. The domain extends axially from -137 cm to 137 cm (i.e., the diameter of the PLX chamber), and spans 68.5 cm radially. The simulation employs adaptive mesh refinement to achieve a maximum effective resolution of nearly 32 cells across the initial outer jet radius.
@@
The jets are modeled as fully-ionized hydrogen with an ideal EOS closure (). The jets are initialized with an inner radius of 1.5 cm (i.e., the plasma gun wire radius) and an outer radius of 4.25 cm. Between these radii, the jets are given a mass density of g cm-3, a velocity of cm s-1, electron and ion temperatures of 1 eV, and a radiation temperature of 0.1 eV. These parameters are enforced as inflow conditions at the lower and upper boundaries of the computational domain, which are also treated as solid reflecting walls from to 1.5 cm and beyond 4.25 cm. The solid-wall regions of the domain boundaries do not allow any hydrodynamic motion, but they allow thermal, radiation, and magnetic diffusion to pass through them. A radially-varying azimuthal magnetic field is imposed for the entire radial extent at these boundaries. That magnetic field has a peak value of 2,500 G at 1.5 cm, with a linear profile from cm to 1.5 cm, a profile from 1.5 cm to 4.25 cm, and a value of zero beyond 4.25 cm. This magnetic field profile is also imposed at every time step on the first row of cells inside the domain, which helps reduce the imbalance of hydromagnetic pressure with the inflow conditions for better control of the jet velocity.
@@
Since Eulerian codes like FLASH do not simulate a true vacuum, the PLX chamber or “ambient” region is initialized with a low density of g cm-3, where thermal conduction and radiation absorption and emission are turned off. The electron and ion temperatures are initialized to 1 eV and radiation temperature to 0.1 eV everywhere inside the domain. The difference in density between the vacuum and jet material creates a pressure imbalance that increases the forward jet velocity and also causes the jets to expand radially into the vacuum. There are also forces from the azimuthal magnetic field that pinch the jet towards during its propagation. Once the pinched jet material reaches the symmetry axis, it has no choice but to move axially, which leads to a streamlining of the jet head, observed as an overall acceleration of the jet; this is a well-documented phenomenon sometimes referred to as the nose-cone effect (see Hansen2015 and references therein). The combined effects of the enforced jet inflow velocity, pressure imbalance, and pinch forces result in an average jet velocity (from s to collision) of about 1.25 cm s-1, which is higher than the reported experimental values Hsu2015 but not necessarily inconsistent. Previous experiments measured the jet velocities at positions near the gun and slightly downstream and reported a possible range of 0.3-1 cm s-1. Given the aforementioned acceleration mechanisms, an average value of 1.25 cm s-1 over the entire axial extent is plausible.
@@
Fig. 1 shows a density and electron temperature map of the jets after they have collided and formed the target plasma. Note that some amount of jet material is left behind or expands outwards away from the target-forming region. This effect can be mitigated by adding more jets; the experimental platform currently employs four target-forming jets, but may use more in the future. Also, some of the jet material will be swept up by the liner-forming jets during the subsequent liner-forming process. As the jets collide, material is forced outwards radially, creating an elongated, pancake-like target. This effect can also be mitigated by adding more jets to the formation process, which would result in a more spherical target. By the end of the simulation, the target reaches a volume-averaged density of approximately g cm-3.
@@
The temperature map of Fig. 1 shows that the hottest plasma forms a shell around the newly formed target. These outermost regions of the target plasma are similar to accretion shocks where jet material piles up and flows into the expanding target. Since the magnetic field is azimuthal and the hot shell has a magnetic resistivity that is lower than that of the jets, the magnetic field has increased difficulty leaving the shell region. The magnetized shell can assist in thermally insulating the target via reduced thermal conductivity, which is included in the anisotropic thermal conduction implementation in FLASH. Some amount of preheat prior to the liner compression of the target is crucial to obtaining the desired plasma properties - we observe peak temperatures of 40 eV in the jet collision simulation.
@@
The postprocessing analysis included volume-averaging several quantities over a region about the center of the domain from to cm and to cm. In this region, the target attains a volume-averaged number density of approximately cm-3, which would increase experimentally with more jets. We also expect the final volume-averaged magnetic field strength, 24.4 G, to also increase with the number of jets. Fig. 2 shows the volume-averaged Hall parameters for electrons () and for ions () as a function of time. The target plasma reaches , which is consistent with previous FLASH and OSIRIS simulations of target formation Wen2022 that reported electron Hall parameters greater than unity. Also, the attained average plasma (i.e., the ratio of thermal-to-magnetic pressure) was slightly greater than 100. This indicates that the magnetic field is strong enough to assist in thermally insulating the target plasma while not being dynamically significant, which is one of the primary goals of the PLX target formation process.
@@
The results from this jet collision simulation are used to inform the initial conditions for 3D experiment-scale target compression simulations presented later in this work (Section V.2).